OpenStudy (anonymous):

how do you work the problem 1 4/5 - 6g = 57

OpenStudy (anonymous):

so it make the equation equal to g you wont to bring all the numbers to the far side so we get -6g=57/(4/5) now when dividing fractions we multiply by there reciprocal . so now we have \[-6g= (57/1) \times (5/4)\] \[-6g = 285/4\] we bring over the -6 \[g = (285/4)/(-6)\] now we again multiply with by the reciprocal \[g= (285/4) \times (1/-6)\] \[g = 285 / -24\] \[g = 11.875\]